OpenStudy (anonymous):

Solve each equation, showing all steps leading to your answer. Then check your answer to determine if any of the solutions are extraneous. 1. Square root sign x+1+5=x 2. Square root sign 5x-x=0 3. Square root sign 2x+1+1=x 4. Square root sign 3x-2+3=4x PLEASE HELP ME!!! :((

OpenStudy (anonymous):

1. \[\sqrt{x+1}+5=x\] 2. \[\sqrt{5x}-x=0\] 3. \[\sqrt{2x+1}+1=x\] 4. \[\sqrt{3x-2}+3=4x\]